OpenStudy (anonymous):

what are some solutions to the equation X^2=49

OpenStudy (anonymous):

please help !!!!!

OpenStudy (anonymous):

7 and -7

OpenStudy (anonymous):

X^2=49 X^2-49=0 (x-7)(x+7)=0 x=7 or x=-7